The perimeter of a rectangle is 420 inches. The width is 4 inches less than the length. What is the length of the rectangle?
Answer by ptaylor(2198)

Let l=length of rectangle
Then l-4=width
perimeter=420=2*length+2*width
420=2l+2(l-4) get rid of parens
420=2l+2l-8 add 8 to both sides
420+8=4l+8-8
428=4l divide both sides by 4
l=107 in-----------------------------length of rectangle
l-4=107-4=103 in---------------------------width of rectangle
CK
2l+2w=2(107)+2(103)=420
214+206=420
420=420

Hope this helps---ptaylor